# How Long Does It Take to Alleviate Discomfort? A Preliminary Study on Reducing Cybersickness in Novice Users

### v1

@@ 1 Introduction
Virtual Reality (VR) is rapidly expanding across domains such as gaming, healthcare, and education, offering users unprecedented immersive interactions [9]. However, cybersickness remains a significant barrier to the widespread adoption of VR technology. Cybersickness occurs when there is a misalignment between visual and vestibular cues and the body’s actual or expected motion [11]. This sensory conflict triggers symptoms such as dizziness and nausea, especially in novice users [11]. From a neurological standpoint, such mismatches are believed to challenge the brain’s ability to reconcile conflicting inputs from the visual and vestibular systems [7]. Although prior research on traditional screen-based media [11], such as 3D games or flight simulators, has shown that rapid scene movements can induce similar discomfort, fully immersive VR tends to exacerbate this mismatch between the visual and vestibular systems.
@@
Existing solutions mainly focus on hardware, software design, and assistive technology, such as improving display resolution, reducing system latency, and adopting dynamic fields of view (FOV) and texture-blurring techniques [4]. Although these methods have proven effective in reducing the incidence and severity of cybersickness, they do not fully address the unique vulnerability of novice users during their initial exposure to VR technology. This heightened susceptibility often leads to discomfort and may discourage prolonged engagement or future use of immersive systems. While system-level technical approaches are valuable, they often overlook individual differences in symptom manifestation and adaptation processes among users [3]. Moreover, the duration and key factors of novice users’ adaptation to immersive environments remain unclear. Additionally, limited research has investigated the progression of cybersickness during prolonged exposure to existing VR applications.
@@
To investigate cybersickness mitigation approaches among novice VR users, we conducted a five-day experiment with 24 undergraduate students with no prior experience in VR. Participants were categorized into mild and high cybersickness groups based on their Day 1 Simulator Sickness Questionnaire (SSQ) [2] scores. Participants joined in VR Jetski Rush and VR C1 Runner within VRChat [8], using tailored levels of locomotion tunneling, a visual mitigation technique based on individual susceptibility. Daily cybersickness levels were measured using the SSQ. Our objectives were to investigate the progression of cybersickness with repeated VR usage and to assess how specific mitigation strategies, such as locomotion tunneling, influence symptom reduction. These findings provide practical guidelines for improving VR accessibility and enhancing user retention, especially among first-time users. By analyzing both quantitative and qualitative data, we gained a deeper understanding of the mechanisms underlying cybersickness among novice VR users.
@@ 2 Related Work
Cybersickness arises from sensory conflict, particularly when visual inputs diverge from vestibular and proprioceptive signals [7]. This mismatch can lead to symptoms such as nausea, dizziness, and fatigue. It is especially prevalent in VR applications that involve rapid movements, sudden scene transitions, or artificially rendered environments, such as flight simulators, 3D games, and amusement rides. Such sensory conflicts are particularly pronounced in scenarios featuring artificial locomotion, which is common in racing or fast-paced games [11]. To mitigate cybersickness, researchers have proposed techniques that reduce the mismatch between visual and inertial cues. These include visual stabilization (e.g., keeping the background static while the foreground moves) [6, 11], image blurring, dynamic FOV restriction [10], and gaze-contingent image deformation to reduce peripheral motion perception [1]. Such strategies aim to harmonize sensory inputs and enhance user comfort in immersive environments. Meanwhile, locomotion tunneling is a technique that reduces users’ peripheral FOV during movement in VR, aiming to minimize visual flow and reduce sensory conflict. A “high locomotion tunneling” condition applies a strong vignette with a narrow central view, while “low locomotion tunneling” offers a mild vignette, retaining more peripheral visibility. In contrast, the “no locomotion tunneling” condition presents the full FOV with no visual restrictions. While system-level technical approaches can partially alleviate cybersickness, individual variability remains a critical factor in the adaptation process. Moreover, there has been limited exploration of how cybersickness evolves over time, particularly among novice VR users.
@@ 3.3 VR Game Selection and Conditions
We used a between-subject design for our experiment. We selected VRChat [8] as the VR experiment platform because it supports remote collaboration and communication among multiple users. It also provided a wide range of virtual entertainment environments. Our goal was to select VR games that are likely to induce cybersickness, such as VR racing games and VR skiing games. Further, we selected VR C1 Runner and VR Jetski Rush. Based on our testing of the two VR games, both exhibited fast-paced visual motion and continuous navigation factors known to increase the risk of cybersickness due to intensified optic flow and sensory conflict [5]. We then confirmed that both VR games supported smooth locomotion, a joystick-based movement mechanism that simulates walking in VR without requiring physical displacement. Moreover, locomotion tunneling was an accessibility feature that included three modes: low, high, and none. It reduced cybersickness for users while playing VR games. Figure 1 A2 & B2 and A3 & B3 show the settings for low and high levels, respectively. Figure 1 A1 and B1 show the original scenes. We deployed VRChat with VR games on two Meta Quest 3 headsets and two Meta Quest 3s headsets.
@@ 3.5 Measurement
We used the SSQ (a 10-point Likert scale) [2] to measure participants’ levels of cybersickness. We also designed a qualitative questionnaire to gather honest feedback from participants, which included six open-ended questions.

### v2

@@ 1 Introduction
Virtual Reality (VR) is rapidly expanding across domains such as gaming, healthcare, and education, offering users unprecedented immersive interactions [WangYi2025]. However, cybersickness remains a significant barrier to the widespread adoption of VR technology. Cybersickness occurs when there is a misalignment between visual and vestibular cues and the body’s actual or expected motion [zhang2016motion]. This sensory conflict triggers symptoms such as dizziness and nausea, especially in novice users [zhang2016motion]. From a neurological standpoint, such mismatches are believed to challenge the brain’s ability to reconcile conflicting inputs from the visual and vestibular systems [oman2014brainstem]. Although prior research on traditional screen-based media [zhang2016motion], such as 3D games or flight simulators, has shown that rapid scene movements can induce similar discomfort, fully immersive VR tends to exacerbate this mismatch between the visual and vestibular systems.
@@
Existing solutions mainly focus on hardware, software design, and assistive technology, such as improving display resolution, reducing system latency, and adopting dynamic fields of view (FOV) and texture-blurring techniques [luks2019investigating]. Although these methods have proven effective in reducing the incidence and severity of cybersickness, they do not fully address the unique vulnerability of novice users during their initial exposure to VR technology. This heightened susceptibility often leads to discomfort and may discourage prolonged engagement or future use of immersive systems. While system-level technical approaches are valuable, they often overlook individual differences in symptom manifestation and adaptation processes among users [kourtesis2024cybersickness]. Moreover, the duration and key factors of novice users’ adaptation to immersive environments remain unclear. Additionally, limited research has investigated the progression of cybersickness during prolonged exposure to existing VR applications.
@@
To investigate cybersickness mitigation approaches among novice VR users, we conducted a five-day experiment with 24 undergraduate students with no prior experience in VR. Participants were categorized into mild and high cybersickness groups based on their Day 1 Simulator Sickness Questionnaire (SSQ) [kennedy1993simulator] scores. Participants joined in VR Jetski Rush and VR C1 Runner within VRChat [VRChat2024], using tailored levels of locomotion tunneling, a visual mitigation technique based on individual susceptibility. Daily cybersickness levels were measured using the SSQ. Our objectives were to investigate the progression of cybersickness with repeated VR usage and to assess how specific mitigation strategies, such as locomotion tunneling, influence symptom reduction. These findings provide practical guidelines for improving VR accessibility and enhancing user retention, especially among first-time users. By analyzing both quantitative and qualitative data, we gained a deeper understanding of the mechanisms underlying cybersickness among novice VR users.
@@ 2 Related Work
Cybersickness arises from sensory conflict, particularly when visual inputs diverge from vestibular and proprioceptive signals [oman2014brainstem]. This mismatch can lead to symptoms such as nausea, dizziness, and fatigue. It is especially prevalent in VR applications that involve rapid movements, sudden scene transitions, or artificially rendered environments, such as flight simulators, 3D games, and amusement rides. Such sensory conflicts are particularly pronounced in scenarios featuring artificial locomotion, which is common in racing or fast-paced games [zhang2016motion]. To mitigate cybersickness, researchers have proposed techniques that reduce the mismatch between visual and inertial cues. These include visual stabilization (e.g., keeping the background static while the foreground moves) [Nie2020, zhang2016motion], image blurring, dynamic FOV restriction [Wu2022AdaptiveFOVRestriction], and gaze-contingent image deformation to reduce peripheral motion perception [Groth2024GazeContingentDeformation]. Such strategies aim to harmonize sensory inputs and enhance user comfort in immersive environments. Meanwhile, locomotion tunneling is a technique that reduces users’ peripheral FOV during movement in VR, aiming to minimize visual flow and reduce sensory conflict. A “high locomotion tunneling” condition applies a strong vignette with a narrow central view, while “low locomotion tunneling” offers a mild vignette, retaining more peripheral visibility. In contrast, the “no locomotion tunneling” condition presents the full FOV with no visual restrictions. While system-level technical approaches can partially alleviate cybersickness, individual variability remains a critical factor in the adaptation process. Moreover, there has been limited exploration of how cybersickness evolves over time, particularly among novice VR users.
@@ 3.3 VR Game Selection and Conditions
We used a between-subject design for our experiment. We selected VRChat [VRChat2024] as the VR experiment platform because it supports remote collaboration and communication among multiple users. It also provided a wide range of virtual entertainment environments. Our goal was to select VR games that are likely to induce cybersickness, such as VR racing games and VR skiing games. Further, we selected VR C1 Runner and VR Jetski Rush. Based on our testing of the two VR games, both exhibited fast-paced visual motion and continuous navigation factors known to increase the risk of cybersickness due to intensified optic flow and sensory conflict [ng2020study]. We then confirmed that both VR games supported smooth locomotion, a joystick-based movement mechanism that simulates walking in VR without requiring physical displacement. Moreover, locomotion tunneling was an accessibility feature that included three modes: low, high, and none. It reduced cybersickness for users while playing VR games. Figure 1 A2 & B2 and A3 & B3 show the settings for low and high levels, respectively. Figure 1 A1 and B1 show the original scenes. We deployed VRChat with VR games on two Meta Quest 3 headsets and two Meta Quest 3s headsets.
@@ 3.5 Measurement
We used the SSQ (a 10-point Likert scale) [kennedy1993simulator] to measure participants’ levels of cybersickness. We also designed a qualitative questionnaire to gather honest feedback from participants, which included six open-ended questions.